# MultiverSeg: Scalable Interactive Segmentation of Biomedical Imaging Datasets with In-Context Guidance

### v1

@@ B.3 Data Augmentation
Appendix C Experimental Setup
C.1 Baselines
@@
C.2 Metrics
@@
Confidence Intervals. We show 95% confidence intervals from bootstrapping. For experiment 1, we bootstrap over results from 200 simulations with different random seeds. For experiment 2, we bootstrap over subjects with 100 runs.
Appendix D Experiment 1: Evaluation
D.1 Setup
@@
Tasks. We exclude tasks with fewer than 18 test examples, leaving 59 tasks from 7 evaluation datasets [2, 11, 3, 1, 40, 118, 142]. We selected this cutoff based on the distribution of task sizes in our validation data (Fig. 9) to focus on scenarios where a user wants to segment many similar images.
@@ D.2 Results
Figure 10: Example segmentation process with MultiverSeg. We begin by interactively segmenting a seed image (Example 0) to 90% Dice. The Example 0 image and final prediction are added to the context set for subsequent examples. For each subsequent example, we first make an initial in-context segmentation prediction using a context set containing all the previous examples and previously predicted segmentations. Then, we simulate center correction clicks until the predicted segmentation achieves $\geq 90\%$ Dice or we have accrued 20 clicks. For Example 2, we only simulated 1 correction because the prediction reached $90\%$ Dice after 1 correction click. For Example 1 and Example 3, additional correction clicks were needed. When the context set is large enough ($>$$n$), the in-context prediction from MultiverSeg may be accurate enough that no corrections are needed. For Example 10, the Dice score of the predicted in-context segmentation is greater than $90\%$ so we do not need to simulate any corrections. In practice, $n$ varies by task.
D.2.1 Interactions per image as a function of dataset size
@@
Results by dataset. As more examples are segmented and the context set grows, the number of clicks and scribbles required to get to $90\%$ Dice on the $\text{n}^{\text{th}}$ example using MultiverSeg decreases substantially. Fig. 11 shows results averaged by dataset. MultiverSeg and SP+UVS are less effective at reducing the number of clicks for tasks from BUID, a breast ultrasound lesion segmentation dataset, perhaps due to the heterogeneity of examples in that dataset.
(a) Center Clicks
(b) Centerline Scribbles
@@
Figure 11: Interactions to target Dice on unseen datasets. Number of interactions needed to reach 90% Dice as a function of the example number being segmented. For the $n^{th}$ image being segmented, the context set has $n$ examples. MultiverSeg requires substantially fewer interactions to achieve 90 Dice than the baselines, and as more images are segmented, the average number of interactions required decreases dramatically. Shaded regions show 95% CI from bootstrapping.
@@
Tasks with more examples. We show results by task for two datasets with more than 18 test examples per task (Fig. 12 and Fig. 13). For larger sets of images, using MultiverSeg results in even greater reductions in the total number of user interactions.
@@
Figure 12: Scribble steps to target Dice by task for BUID. Number of interactions needed to reach a 90% Dice as a function of the example number being segmented. For the $n^{th}$ image being segmented, the context set has $n$ examples. Shading shows 95% CI from bootstrapping. BUID [3] is a breast ultrasound dataset containing segmentation tasks for benign and malignant lesions. As the context set of completed segmentations grows, the number of interactions required to segment each additional image with MultiverSeg declines.
@@
Context Set Quality. For MultiverSeg and SP+UVS, thresholding the predictions before adding them to the context set improved performance (Fig. 14). We use the validation split of our validation data to select the best approach (soft or binary predictions in the context set) for each method.
@@
SP+UVS. Consistent with the original published results, we find that UniverSeg has poor performance for small context sets and initializing ScribblePrompt using the UniverSeg prediction hurts performance when the context set is small. In our final evaluation of SP+UVS, we set the minimum context set size to be 5 examples: when the context sets contains fewer than 5 examples, we ignore the context and only use ScribblePrompt to make predictions. Fig. 15 shows variations of SP+UVS with different minimum context set sizes.
@@
Total Interactions. Fig. 16 shows the total number of interaction steps per set of 18 images by dataset. Tab. 7 shows the total number of interactions, average Dice score, and average 95th percentile Hausdorff distance across all tasks.
Figure 16: Total interactions per unseen task. We show total number of clicks and scribble steps to segment 18 images to $\geq 90\%$ Dice for each method. In all scenarios, MultiverSeg required fewer interactions than both baselines. Error bars show 95% CI from bootstrapping.
@@
Center Clicks ScribblePrompt $91.23\pm 0.08$ $1.04\pm 0.02$ $131.03\pm 1.32$
SP+UVS $91.14\pm 0.11$ $1.07\pm 0.03$ $115.30\pm 2.34$
MultiverSeg (ours) $\mathbf{91.35\pm 0.21}$ $\mathbf{1.00\pm 0.04}$ $\mathbf{84.39\pm 2.58}$
Centerline Scribbles ScribblePrompt $\mathbf{93.23\pm 0.21}$ $0.90\pm 0.10$ $55.27\pm 3.28$
SP+UVS $93.22\pm 0.20$ $0.90\pm 0.12$ $42.77\pm 2.75$
MultiverSeg (ours) $\mathbf{93.23\pm 0.21}$ $\mathbf{0.86\pm 0.10}$ $\mathbf{26.13\pm 2.39}$
Table 7: Average segmentation quality and total interactions per unseen task. We measure average segmentation quality across a set of 18 test images using Dice score and 95th percentile Hausdorff distance (HD95). For each metric, we show mean and standard deviation from bootstrapping. Dice and HD95 are similar across methods because we simulate interactions until the predicted segmentation has $\geq 90\%$ Dice or the maximum number of interaction steps is reached. MultiverSeg requires the fewest interaction steps per task on average.
D.2.2 Bootstrapping In-Context Segmentation
@@
UniverSeg $51.02\pm 0.07$ $16.60\pm 0.44$
UniverSeg (oracle) $70.24\pm 0.13$ $13.28\pm 0.28$
@@
Appendix E Experiment 2: Analysis
E.1 In-Context Segmentation
@@
Figure 19: In-context segmentation performance across context set sizes on unseen datasets. We compare MultiverSeg to UniverSeg, an in-context segmentation method, given ground truth context labels. Shading shows 95% CI from bootstrapping.
E.2 Interactive Segmentation In Context

### v2

@@ C.3 Data Augmentation
Appendix D Experimental Setup
D.1 Baselines
@@
SAM [61] $\checkmark$
MedSAM [88] $\checkmark$
SAM-Med2D [19] $\checkmark$
SegNext [79] $\checkmark$
@@
SAM. We evaluated SAM [61] (ViT-b) in both “single-mask” and “multi-mask” mode on our validation data, and average results were better using “single-mask” mode. We report final results for SAM on the test data using “single-mask” mode.
@@
D.2 Inference
Image Resolution. MultiverSeg, ScribblePrompt, and UniverSeg, which were all developed and trained on $128^{2}$ sized images, and output predictions at the same resolution. SAM was trained with $1024^{2}$ sized inputs and predicts segmentations at $256^{2}$ resolution. For each method, we resized the inputs to the method’s training input size using bilinear interpolation before performing inference and then resized the output (as needed) to the evaluation resolution.
D.3 Metrics
@@
Confidence Intervals. For Experiment 1, we calculate 95% confidence intervals over results from 200 simulations with different random seeds. For Experiment 2, we calculate 95% confidence intervals by bootstrapping over subjects with 100 runs.
Appendix E Experiment 1: Evaluation
E.1 Setup
@@
Tasks. We exclude tasks with fewer than 18 test examples, leaving 161 tasks from 8 evaluation datasets [2, 11, 3, 1, 42, 122, 146, 129]. We selected this cutoff based on the distribution of task sizes in our validation data (Fig. 9) to focus on scenarios where a user wants to segment many similar images.
Data. We conducted our evaluation on $256^{2}$ sized images. For each method, we resized the inputs to match the size of the model’s training data before performing the forward pass, and then resized the prediction back to $256^{2}$ before calculating the Dice Score. In Sec. E.5 we conduct a sensitivity analysis, performing the evaluation with $128^{2}$ sized images
@@
Figure 10: Example segmentation process with MultiverSeg. We begin by interactively segmenting a seed image (Example 0) to 90% Dice. The Example 0 image and final prediction are added to the context set for subsequent examples. For each subsequent example, we first make an initial in-context segmentation prediction using a context set containing all the previous examples and previously predicted segmentations. Then, we simulate center correction clicks until the predicted segmentation achieves $\geq 90\%$ Dice or we have accrued 20 clicks. For Example 2, we only simulated 1 correction because the prediction reached $90\%$ Dice after 1 correction click. For Example 1 and Example 3, additional correction clicks were needed. When the context set is large enough (>>$n$), the in-context prediction from MultiverSeg may be accurate enough that no corrections are needed. For Example 10, the Dice score of the predicted in-context segmentation is greater than $90\%$ so we do not need to simulate any corrections. In practice, $n$ varies by task.
E.2 Interactions per Image as a Function of Dataset Size
@@
Results by dataset. As more examples are segmented and the context set grows, the number of clicks and scribbles required to get to $90\%$ Dice on the $\text{n}^{\text{th}}$ example using MultiverSeg decreases substantially. Fig. 11 and Fig. 12 show results averaged by dataset. MultiverSeg and SP+UVS are less effective at reducing the number of clicks for tasks from BUID, a breast ultrasound lesion segmentation dataset, perhaps due to the heterogeneity of examples in that dataset.
@@
Figure 11: Clicks to target Dice on unseen datasets. Number of interactions needed to reach 90% Dice as a function of the example number being segmented. For the $n^{th}$ image being segmented, the context set has $n$ examples. MultiverSeg requires substantially fewer interactions to achieve 90 Dice than the baselines, and as more images are segmented, the average number of interactions required decreases dramatically. Shaded regions show 95% CI from bootstrapping.
Figure 12: Scribbles to target Dice on unseen datasets. Number of interactions needed to reach 90% Dice as a function of the example number being segmented. For the $n^{th}$ image being segmented, the context set has $n$ examples. MultiverSeg requires substantially fewer interactions to achieve 90 Dice than the baselines, and as more images are segmented, the average number of interactions required decreases dramatically. Shaded regions show 95% CI from bootstrapping.
@@
Tasks with more examples. We show results by task for three datasets with more than 18 test examples per task (Fig. 13, Fig. 14, and Fig. 15). For larger sets of images, using MultiverSeg results in even greater reductions in the total and average number of user interactions.
@@
Figure 13: Scribble steps to target Dice by task for WBC. Number of interactions needed to reach a 90% Dice as a function of the example number being segmented. For the $n^{th}$ image being segmented, the context set has $n$ examples. Shading shows 95% CI from bootstrapping. WBC [146] is a microscopy dataset containing segmentation tasks for cytoplasm and nuclei of white blood cells. After segmenting a few images from the femur task with MultiverSeg, the rest of the images in the task can be segmented (to $\geq 90\%$ Dice) with minimal (or no) additional interactions.
Figure 14: Scribble steps to target Dice by task for BUID. Number of interactions needed to reach a 90% Dice as a function of the example number being segmented. For the $n^{th}$ image being segmented, the context set has $n$ examples. Shading shows 95% CI from bootstrapping. BUID [3] is a breast ultrasound dataset containing segmentation tasks for benign and malignant lesions. As the context set of completed segmentations grows, the number of interactions required to segment each additional image with MultiverSeg gradually declines.
@@
Context Set Quality. For MultiverSeg and SP+UVS, thresholding the predictions before adding them to the context set improved performance (Fig. 16). We use the validation split of our validation data (at $128^{2}$ resolution) to select the best approach (soft or binary predictions in the context set) for each method.
@@
SP+UVS. Consistent with the original published results, we find that UniverSeg has poor performance for small context sets and initializing ScribblePrompt using the UniverSeg prediction hurts performance when the context set is small. In our final evaluation of SP+UVS, we set the minimum context set size to be 5 examples: when the context sets contains fewer than 5 examples, we ignore the context and only use ScribblePrompt to make predictions. Fig. 17 shows variations of SP+UVS with different minimum context set sizes on validation data at $128^{2}$ resolution.
@@
Total Interactions. Fig. 18 shows the total number of interactions, average Dice score, and average 95th percentile Hausdorff distance across all tasks.
@@
Center Clicks SAM-Med2D $85.88\pm 0.14$ $3.76\pm 0.22$ $215.58\pm 2.22$
IMIS-Net $81.38\pm 0.30$ $13.05\pm 0.79$ $255.47\pm 2.53$
SAM $90.40\pm 0.06$ $1.40\pm 0.03$ $152.55\pm 1.76$
SegNext $90.50\pm 0.05$ $1.84\pm 0.06$ $158.16\pm 0.95$
ScribblePrompt $90.80\pm 0.08$ $1.48\pm 0.04$ $137.10\pm 1.21$
SP+UVS $90.70\pm 0.09$ $1.49\pm 0.06$ $122.01\pm 1.93$
MultiverSeg (ours) $\mathbf{91.40\pm 0.14}$ $\mathbf{1.26\pm 0.11}$ $\mathbf{87.18\pm 1.92}$
Centerline Scribbles SAM-Med2D $29.58\pm 3.92$ $26.42\pm 3.36$ $178.00\pm 1.19$
IMIS-Net $80.93\pm 0.40$ $3.43\pm 0.32$ $123.46\pm 2.85$
SAM $80.19\pm 0.74$ $19.79\pm 1.78$ $125.14\pm 2.56$
ScribblePrompt $88.19\pm 0.24$ $\mathbf{1.44\pm 0.06}$ $100.70\pm 2.67$
SP+UVS $88.57\pm 0.23$ $1.44\pm 0.07$ $92.50\pm 1.95$
MultiverSeg (ours) $\mathbf{88.65\pm 0.22}$ $1.49\pm 0.13$ $\mathbf{75.23\pm 1.50}$
Figure 18: Average segmentation quality and total interactions per unseen task. We measure average segmentation quality across a set of 18 test images using Dice score and 95th percentile Hausdorff distance (HD95). For each metric, we show mean and standard deviation from bootstrapping. Dice and HD95 are similar across methods because we simulate interactions until the predicted segmentation has $\geq 90\%$ Dice or the maximum number of interaction steps is reached. MultiverSeg requires the fewest interaction steps per task on average. We report results on images at $256^{2}$ resolution from 200 simulations.
E.3 Bootstrapping In-Context Segmentation
@@
UniverSeg $48.89\pm 1.87$ $16.76\pm 0.40$
UniverSeg (oracle) $68.15\pm 1.00$ $13.58\pm 0.24$
@@
E.4 Comparison to Few-Shot Fine-Tuning
One approach to segmenting a new dataset is to (interactively) segment a few images using a pre-trained foundation model, and then use those examples to train a task-specific interactive segmentation model by fine-tuning the foundation model. In this experiment, we simulated this process using ScribblePrompt.
Setup. For each task and random seed, we sampled 5 random test examples, and used ScribblePrompt to segment those images using simulated random center clicks. For each image of the 5 images, random center clicks were used to prompt ScribblePrompt until a maximum of 20 clicks was reached or the prediction surpassed 90% Dice. Then we used those newly labeled images to fine-tune ScribblePrompt from pre-trained weights. We randomly split the 5 images into 4 training examples and 1 validation example.
We fine-tuned ScribblePrompt using the same training interaction protocol, loss function, and data augmentations (Sec. C.3) as MultiverSeg minus synthetic task augmentations. Each task-specific model was fine-tuned for 300 epochs using the Adam optimizer with a learning rate of $1e^{-6}$ and batch size of $4$. These hyperparameters were selected based on experiments with learning rate $\in\{1e^{-4},1e^{-5},1e^{-6}\}$ and batch size $\in\{4,8\}$ using the cytoplasm segmentation task from the WBC [146] dataset. For each training run the best checkpoint was selected based on the validation example and then used to interactively segment 13 more test images (to complete the set of 18).
We repreated this procedure of labelling images and training tasks-specific models for 5 random seeds for each task. Due to the large number of tasks-specific models trained for this experiment, we trained and evaluated on images at $128^{2}$ to reduce training time.
Runtime. Fine-tuning ScribblePrompt to produce each task-specific interactive segmentation model took on average 20 minutes on a NVIDIA A100 GPU. In contrast, MultiverSeg’s inference time is <150<150 milliseconds, even with a context set size of 64 examples (Sec. F.3).
Results. Fig. 21 shows MultiverSeg required fewer interactions than fine-tuning ScribblePrompt in 13 out of 16 scenarios. On average, the fine-tuning approach required $5.90\pm 0.10$ clicks or $2.63\pm 0.13$ scribble steps per image. MultiverSeg required fewer interactions: $4.64\pm 0.10$ clicks or $4.64\pm 0.10$ scribble steps per image.
Figure 21: MultiverSeg outperforms task-specific fine-tuning on most datasets. We show average number of clicks and scribble steps per image to segment 18 images to $\geq 90\%$ Dice for each method. For FT ScribblePrompt (shaded), we used ScribblePrompt to interactively segment 5 images and then used those examples to fine-tune ScribblePrompt before interactively segmenting the rest. MultiverSeg required fewer interactions thant fine-tuned ScribblePrompt in 13 out of 14 scenarios. Error bars show 95% CI accross 5 random seeds.
E.5 Resolution Sensitivity Analysis
We conduct a sensitivity analysis, evaluating MultiverSeg and the baseline methods at $128^{2}$ resolution.
Results. MultiverSeg outperforms the baselines with greater margins when evaluated at $128^{2}$ resolution compared to $256^{2}$ resolution. As more examples are segmented and the context set grows, the number of interactions required to get to $90\%$ Dice (NoI90) on the $\text{n}^{\text{th}}$ example using MultiverSeg decreases substantially (Fig. 22).
MultiverSeg required the fewest number of interactions per image on all datasets (Fig. 23). On average, using MultiverSeg reduced the number of clicks required to segment each dataset by $(36.93\pm 1.53)\%$ and the number of scribble steps required by $(36.93\pm 1.53)\%$ compared to ScribblePrompt.
Figure 22: Interactions to target Dice on unseen tasks at $\mathbf{128^{2}}$ resolution. Number of interactions needed to reach a 90% Dice as a function of the example number being segmented. For the $n^{th}$ image being segmented, the context set has $n$ examples. MultiverSeg requires substantially fewer number of interactions to achieve 90% Dice than the baselines, and as more images are segmented, the average number of interactions required decreases dramatically. Shaded regions show 95% CI accross 200 random seeds.
Figure 23: Interactions per image by unseen dataset at $\mathbf{128^{2}}$ resolution. We show average number of clicks and scribble steps per image to segment 18 images to $\geq 90\%$ Dice for each method. In all scenarios, MultiverSeg required fewer or the same number of interactions than the best baseline. Error bars show 95% CI accross 200 random seed.
Appendix F Experiment 2: Analysis
F.1 In-Context Segmentation
@@
Figure 24: In-context segmentation performance across context set sizes on unseen datasets. We compare MultiverSeg to UniverSeg, an in-context segmentation method, given ground truth context labels. Points show results for context set sizes 1, 2, 4, 8, 16, 32, 64, 96, 128 and 256. Shading shows 95% CI from bootstrapping.
F.2 Interactive Segmentation In Context
@@
F.3 Inference Runtime and Memory Usage
MultiverSeg’s inference runtime scales linearly with the context set size (Tab. 7). However, even with a context set of 64 examples, the runtime is under 150150ms. Prior work on interactive interfaces indicates <500<500ms latency is sufficient for cognitive tasks [80]. Since the interactions are stored in masks, inference runtime (per prediction) is not affected by the number of user interaction inputs.
Context Size Inference Time (ms) GPU Memory
1 $25.28\pm 0.16$ 28 MB
16 $57.05\pm 0.20$ 1.89 GB
32 $86.57\pm 0.06$ 3.64 GB
64 $146.04\pm 0.16$ 7.15 GB
128 $267.42\pm 0.24$ 12.16 GB
256 $604.15\pm 0.36$ 24.17 GB
Table 7: Inference runtime and GPU memory usage with different context set (CS) sizes. We report mean $\pm$ standard deviation runtime in milliseconds across 1,000 predictions at $128^{2}$ resolution with 1 click on an NVIDIA A100 GPU. GPU memory usage is reported as peak allocated memory during inference.